Determine the solubility of carbon dioxide in water at 25 degrees C exposed to air at 1.0 atm. Assume a partial pressure for carbon dioxide of 9.3x10^-3 atm. (kH,CO2=3.4×10−2M/atm.)
Answer #1

Solution :-

Using the Henrys law we can calculate the solubility of the gas

C= kH* Pgas

C= 3.4*10^-2 M/atm * 9.3*10^-3 atm

C= 3.162*10^-4 M

Therefore the solubility of the CO2 gas is

3.162*10^-4 mol per L * 44.01 g per mol =0.0139 g/L